# On the Design and Analysis of LLM-Based Algorithms

### v2

@@ Abstract
We initiate a formal investigation into the design and analysis of LLM-based algorithms, i.e. algorithms that contain one or multiple calls of large language models (LLMs) as sub-routines and critically rely on the capabilities of LLMs. While LLM-based algorithms, ranging from basic LLM calls with prompt engineering to complicated LLM-powered agent systems and compound AI systems, have achieved remarkable empirical success, the design and optimization of them have mostly relied on heuristics and trial-and-errors, which is largely due to a lack of formal and analytical study for these algorithms. To fill this gap, we start by identifying the computational-graph representation of LLM-based algorithms, the design principle of task decomposition, and some key abstractions, which then facilitate our formal analysis for the accuracy and efficiency of LLM-based algorithms, despite the black-box nature of LLMs. Through extensive analytical and empirical investigation in a series of case studies, we demonstrate that the proposed framework is broadly applicable to a wide range of scenarios and diverse patterns of LLM-based algorithms, such as parallel, hierarchical and recursive task decomposition. Our proposed framework holds promise for advancing LLM-based algorithms, by revealing the reasons behind curious empirical phenomena, guiding the choices of hyperparameters, predicting the empirical performance of algorithms, and inspiring new algorithm design. To promote further study of LLM-based algorithms, we release our source code at https://github.com/modelscope/agentscope/tree/main/examples/paper_llm_based_algorithm.
@@ 1 Introduction
The rapid advancements of pre-trained large language models (LLMs) in the past few years [BCE+23, WTB+22, SMK23] have given rise to the new paradigm of algorithmic problem solving by utilizing LLMs as general-purpose solvers and prompting them with task descriptions plus additional inputs that can help enhance their performance [WWS+22, KGR+22, ZYY24]. Meanwhile, it is also well recognized that even the best LLMs today still exhibit various limitations, such as finite context window sizes and difficulty with complex reasoning. Some of these limitations are fundamental [MS23, PNP24, TTC+24, HR24, WDL24, MPS24], and resolving them requires new breakthroughs. Moreover, in resource-constrained scenarios where using the state-of-the-art LLMs is not feasible, one might have to resort to smaller and weaker LLMs for solving complex tasks.
@@
All these have motivated the developments of what we call LLM-based algorithms, namely, algorithms that contain one or multiple LLM calls as sub-routines and fundamentally rely on the capabilities of LLMs. In its most basic form, an LLM-based algorithm can be a combination of one or multiple LLM calls and some prompt engineering [YYZ+23, BBK+23, ZSH+23, WWS+23]. More advanced examples include LLM-powered agent systems [MDL+23, PKB+24, XCG+23] and compound AI systems [ZKC+24] that augment LLMs with additional abilities like tool use and long-term memory, as well as the emerging paradigm of LLM programming [SSC+23, KSM+24, ZYX+24].
LLM-based algorithms, in a similar spirit to neuro-symbolic programming [CEP+21, GK23] and learning-augmented algorithms [MV22, LM22], combine the advantages of both LLMs and traditional algorithms. An LLM-based algorithm, designed with human’s intelligence and knowledge of algorithmic problem solving, can exhibit much better controllability and interpretability, stronger performance that is less reliant on delicate prompting or extensive trial-and-errors, and capabilities that far exceed what could possibly be achieved by directly prompting the LLM for a solution.
The rapid developments of LLM-based algorithms naturally raise the question: is it possible to provide any formal analysis or guarantee for LLM-based algorithms? This seems like a daunting task at first glance, due to the black-box nature of LLMs. Indeed, in prior works of this field, LLM-based algorithms have been mostly designed in a heuristic manner, and evaluated empirically with limited insights beyond the measurements of a few error or cost metrics on certain benchmarks. In contrast, formal analysis of LLM-based algorithms, if it does exist, can bring various potential benefits, including but not limited to revealing the reasons behind curious empirical phenomena, instructing choices of hyperparameters, predicting the empirical performance of LLM-based algorithms, and even inspiring new algorithm design.
@@
The goal of this work is to initiate a formal investigation into the design and analysis of LLM-based algorithms. Our contributions to filling this gap, and thereby advancing the field of LLM-based algorithms, are summarized as follows.
  • In Section 2, we start by formulating LLM-based algorithms as computational graphs consisting of LLM nodes and non-LLM nodes, and identifying the design principle of task decomposition. We further introduce some key abstractions, based on which formal analysis for the accuracy and efficiency of LLM-based algorithms is developed.
  • Section 3 provides an in-depth case study for parallel decomposition, a basic pattern of task decomposition and a building block for more sophisticated LLM-based algorithms. After introducing the general formulation and analysis for this pattern, we present our algorithm design and analysis for multiple concrete tasks, which are then validated by extensive numerical experiments. Advanced case studies for hierarchical decomposition and recursive decomposition are provided in Sections 4 and 5 respectively, which further demonstrate how to apply our proposed analytical framework to diverse scenarios and complex patterns of LLM-based algorithms.
  • We derive various novel insights from our analytical and empirical study. For example, considering the hyperparameter $m$ that represents the granularity of parallel decomposition, our analysis explains why error and cost metrics of an LLM-based algorithm are monotone in $m$ in certain cases while non-monotone in others, which in turn guides the choices of $m$ for achieving the desired accuracy or efficiency. Our work exemplifies how to leverage the proposed framework for systematic design and analysis of practical LLM-based algorithms, which can potentially inspire future work in this area.
2 LLM-based algorithms
This section introduces a formal framework for the design and analysis of LLM-based algorithms. In the following, we present an overview for the definition, design principle, and analysis of LLM-based algorithms. Further details are provided afterwards.
Definition.
Generally speaking, an LLM-based algorithm is simply an algorithm that contains one or multiple LLM calls as its key components. Examples of LLM-based algorithms range from one single LLM call, to LLM-powered agent systems or compound AI systems consisting of a mixture of LLM calls and non-LLM programs. One way of formally representing an LLM-based algorithm is using a computational graph, which will soon be elaborated in Section 2.1.
Design principle: task decomposition.
An LLM-based algorithm might utilize multiple LLM calls and non-LLM programs to handle sub-tasks derived from the original problem, whose outputs together give rise to the final solution of the overall algorithm. This naturally implies the principle of task decomposition. Indeed, one key aspect of designing an LLM-based algorithm is to figure out how to decompose the original task into sub-tasks appropriately, so that each of them can be handled by one LLM call or non-LLM program accurately and efficiently, while the performance of the overall algorithm is also guaranteed.
The principle of task decomposition has been widely adopted, either explicitly or implicitly, in prior works on LLM-based algorithms. There are various reasons why task decomposition is beneficial and oftentimes crucial for solving problems with LLMs, some of which are listed below:
  • Despite the rapid developments of Transformer-based LLMs, their context windows are still finite, which limits the problem size that one LLM call can handle. For an LLM that have a very large or even (in theory) infinite context window, its accuracy of solving a class of problems, e.g., long-text retrieval or summarization, will typically decay with the size of the problem, in which case task decomposition is still favorable.
  • Other classes of problems might have limited problem sizes, but require complex reasoning for solving them correctly. In this case, LLM-based algorithms can benefit substantially from sequential task decomposition such as step-by-step thinking and acting [YZY+23], or parallel decomposition like generating multiple reasoning paths and then aggregating them for a better solution [SLC+23, WWS+23].
  • Aside from accuracy, it is possible that better efficiency can be achieved with task decomposition, compared with one single LLM call. For instance, the input problem might be decomposed into multiple independent sub-tasks that can be solved in parallel, while solving the problem with one LLM call via autoregressive decoding, which is inherently sequential, suffers from a larger end-to-end latency.
Despite all these benefits, it is also possible that fine-grained task decomposition can incur higher errors or costs in certain cases. Choosing the appropriate decomposition is crucial for achieving good performance of the overall algorithm, be it a trade-off between accuracy and efficiency or the best of both worlds.
Remark 1.
It is worth noting that one single LLM call can also be treated from the perspective of task decomposition, which is a special case of LLM-based algorithms. Thus the analysis proposed in this work still holds for this case.
Analysis: accuracy and efficiency.
Given a task and the corresponding LLM-based algorithm with specific configurations, we aim to analyze its performance, namely how accurately and efficiently the algorithm solves the task, akin to analysis for any generic algorithm. Following the principle of task decomposition, this is done by analyzing for each LLM call or non-LLM program first, and then for the overall algorithm. Given that LLMs are regarded as black-box general-purpose problem solvers using natural language as input and output, we find it useful and necessary to leverage certain abstractions, to be introduced in Section 2.2, in order to facilitate formal analysis that will soon be presented in Section 2.3. One practical usage of formal analysis is to predict the performance of an LLM-based algorithm before actually running it, which will in turn help to optimize certain hyperparameters in the configurations or inspire better algorithm design.
2.1 Algorithms as computational graphs
We formulate an LLM-based algorithm as a computational graph. Each graph node takes some inputs from its predecessor nodes, executes certain operations, and returns some outputs. The nodes can be categorized into two types, which we refer to as LLM nodes and non-LLM nodes, demonstrated in Figure 1.
@@ 2.1 Algorithms as computational graphs
  • Within an LLM node, the operations consist of a prompter that formats a prompt based on the inputs, an LLM call that processes the prompt, and a parser that extracts the targeted information from the LLM’s response. The prompter and parser, designed for the sub-task of the current node, serve as translators between natural language and traditional data structures. Following the principle of task decomposition, we assume without loss of generality that each LLM node contains one single LLM call.
  • Within a non-LLM node, the operations can be anything that does not involve LLMs. Examples include a symbolic algorithm, an API call for a search engine, or a classical machine learning model.
@@
Given such nodes, the computational graph is built by connecting them with directed edges that specify the data flows within the LLM-based algorithm. For example, Figure 2(a) shows the graph for the pattern of parallel decomposition, which divides the input problem into parallel sub-tasks, solves each of them with one LLM node, and aggregates the results for the final solution; Figure 2(b) illustrates an algorithm for book-length summarization [CLGI24, Ope24b], which divides the input text into multiple chunks and maintains a global summary via incremental updating; and Figure 2(c) represents the ReAct algorithm [YZY+23], which consists of multiple iterations of reasoning, tool use and aggregation. The computational-graph formulation is expressive enough to cover these diverse LLM-based algorithms that have been proposed separately in prior works, and facilitates formal analysis for all of them in a unified manner. The computational graph of an LLM-based algorithm can be pre-specified and static, or dynamically constructed at runtime, as will be demonstrated later in our case studies in Sections 4 and 5. Configurations of an LLM-based algorithm include how task decomposition is done (reflected by the graph topology and the sub-tasks of graph nodes), the methods of prompting and parsing for each LLM node, configurations of the LLM(s) being used, and so on. The LLM nodes within one graph might use the same or different backbone LLMs.
@@
Figure 1: Two types of nodes in the computational graphs of LLM-based algorithms. Each node can have one or multiple inputs/outputs. We use the abbreviation “NL” for “natural language”, and “DS” for “data structure”.
(a) Parallel decomposition. See Section 3 for four concrete examples of this pattern.
(b) Book-length summarization, cf. Figure 1 in [CLGI24]. The “dividing” node contains a symbolic program that divides the input text into multiple smaller chunks.
(c) The ReAct algorithm [YZY+23]. Each “acting” node represents one API call for some tool, and each “aggregation” node aggregates the outputs of its predecessor nodes, e.g. by simple concatenation.
Figure 2: Examples of computational-graph representations for LLM-based algorithms.
2.2 Key abstractions
Let us introduce a few key abstractions that facilitate our formal analysis of LLM-based algorithms, namely, how accurately and efficiently an LLM-based algorithm solve a task. A summary of these abstractions and their relation can be found in Figure 3.
Figure 3: The introduced abstractions and their relation.
@@ 2.2 Key abstractions
Error metrics are task-specific in general. Moreover, the overall algorithm and the LLM nodes within it might have different error metrics. For instance, consider an LLM-based algorithm that answers a given question by first retrieving relevant sentences from the input text with multiple LLM nodes, and then reasoning over the extracted sentences with another LLM node. In this case, the overall algorithm is evaluated by whether it answers the question correctly, while each LLM node responsible for retrieval is evaluated by whether it successfully extracts all necessary sentences within the input text that it handles.111 Throughout this work, we use “accuracy” to refer to the broader concept of “quality”, and an “error metric” can be any metric that measures how much the output of an algorithm deviates from certain criteria.
Cost metrics, on the other hand, are largely task-agnostic. Below are some example cost metrics that might be of interest, regardless of the specific task or algorithm:
  • The total length of prompts and total length of generated texts, measured by the numbers of characters or tokens, within one run of the overall algorithm. These metrics are especially relevant to the financial costs when the algorithm uses a proprietary LLM that is accessed via commercial API calls, which are typically charged by the number of tokens. They are also correlated with the time and memory complexities of LLM inference.
  • The end-to-end latency of one run of the overall algorithm, namely the time complexity. One particular property of this metric is that the latency can be impacted by parallelism of LLM calls, an important aspect of LLM inference service in practice.
  • Other possible cost metrics for one run of an LLM-based algorithm include the peak memory usage, the total number of LLM calls, FLOPs, energy consumption, carbon emission, among others.
Unless specified otherwise, we focus on the costs of the LLM nodes within an LLM-based algorithm, and neglect those of the non-LLM nodes, since the latter is much smaller than the former in all concrete scenarios that we will consider later in this work.
@@
To analyze the error and cost metrics for an LLM-based algorithm, it is necessary to make certain assumptions about the characteristics and inference service of the LLM(s) used within it.
@@
Assumptions on LLM inference service [YSZ+24, PDC+23, ZNH+24], on the other hand, are task-agnostic and only affect the cost metrics, not error metrics. In particular, they determine the relation between the cost metrics and the lengths of prompt and generated text for each LLM call, the parallelism of multiple LLM calls, among others. While LLM inference service in practice can be very diverse, we will see in the next subsection that unified and formal analysis is possible with appropriate abstractions of it.
2.3 Formal analysis: accuracy and efficiency
Given the aforementioned formulations and abstractions, we are now ready to consider formal analysis for the accuracy and efficiency of LLM-based algorithms. More specifically, one can first analyze the error and cost metrics for each individual node within the computational graph; these, combined with the graph topology, lead to results about the error and cost metrics of the overall algorithm.
We elaborate this approach in the following. Unless otherwise specified, our analysis is deterministic, for a given task instance and fixed random seed(s).
Analysis of error metrics.
The error metrics of the output of each LLM node, which are calculated with respect to what the output should have been if all nodes accomplish their tasks with exact accuracy, depend on the characteristics, i.e. capabilities and limitations, of the LLM, as well as the specific problem instance and random seed(s). For a certain node $v$, the error of its output $y$ can be bounded by some function $f_{v}$ (which depends on the aforementioned factors if $v$ is an LLM node) of the errors of its inputs $x_{1},x_{2},\dots,x_{k}$, i.e. the outputs of its predecessor nodes:
@@ 2.3 Formal analysis: accuracy and efficiency
The function $f_{v}$ can be general, with some examples listed in the following:
  • $\mathcal{E}(y)\leq\sum_{i\in[k]}\mathcal{E}(x_{i})$ or $\frac{1}{k}\sum_{i\in[k]}\mathcal{E}(x_{i})$. Such linear relationship can appear in, for example, a simple counting task that will be elaborated in Section 3.3.
  • $\mathcal{E}(y)=\min_{i\in[k]}\mathcal{E}(x_{i})$. In a task of code generation, the algorithm might generate multiple samples, so that it succeeds as long as one of the samples is correct (e.g. by passing all test cases).
  • $\mathcal{E}(y)\leq\max_{i\in[k]}\mathcal{E}(x_{i})$. This case can be used for the sorting task to be introduced in Section 3.4, if the node $v$ is responsible for merging multiple sorted lists and $\mathcal{E}$ stands for the $\ell_{\infty}$ error.
Finally, the error metrics of the overall algorithm are exactly those of the particular graph node that is responsible for generating the final solution returned by the algorithm.
Analysis of cost metrics.
@@
  • Charges of API calls: $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$ are linear functions, namely $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})=c_{\textsf{pre}}\times\mathcal{L}_{\textsf{pre}}$, $\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})=c_{\textsf{dec}}\times\mathcal{L}_{\textsf{dec}}$, where $c_{\textsf{pre}}$ and $c_{\textsf{dec}}$ are the cost per token for prompting and text generation, respectively.
  • Latency: in a memory-bound scenario [AQS+23] where $\mathcal{L}_{\textsf{pre}}$ is limited and the latency of memory IO for loading model weights is dominant, we have approximately $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})=O(1)$ and $\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})=O(\mathcal{L}_{\textsf{dec}})$; in more general scenarios, $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$ might grow linearly or even quadratically (e.g. for Transformers with full attention [VSP+17]) with $\mathcal{L}_{\textsf{pre}}$ and $\mathcal{L}_{\textsf{dec}}$.
Next, we consider the cost of the overall algorithm, which can be written as a function of the cost metrics of all LLM nodes within its computational graph. It might be a simple sum of the costs of all LLM nodes, e.g. for API costs charged by tokens. Or it can be a more complex function, e.g. for the end-to-end latency in the presence of parallelism. As an example, consider $k$ independent LLM calls with latencies $\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{k}$ and ideal parallelism of degree $p$, i.e. no extra cost is induced by parallelism; then, the $k$ LLM calls can be divided into $g=\lceil k/p\rceil$ sequential groups, where each group of $p$ parallel LLM calls is bottlenecked by the one with the largest latency, and thus the end-to-end latency of the overall algorithm becomes
@@
2.4 Practical considerations
Let us comment on a few practical considerations about the aforementioned abstractions and analysis.
  • Regarding the capabilities and limitations of LLMs in a specific task, the black-box nature of LLMs can make it challenging to analytically and accurately quantify these factors, in which case one might resort to measuring and profiling in practice. On the positive side, it is oftentimes easier to make certain qualitative assumptions. For example, in many tasks of interest, a larger problem instance is harder than a smaller one, and thus incurs larger error metrics of an LLM call. For practical purposes like optimizing certain hyperparameters of an LLM-based algorithm, such weak assumptions might be sufficient already.

### v3

# Designing Algorithms Empowered by Language Models: An Analytical Framework, Case Studies, and Insights

@@ Abstract
This work presents an analytical framework for the design and analysis of LLM-based algorithms, i.e., algorithms that contain one or multiple calls of large language models (LLMs) as sub-routines and critically rely on the capabilities of LLMs. While such algorithms, ranging from basic LLM calls with prompt engineering to complicated LLM-powered agentic workflows and compound AI systems, have achieved remarkable empirical success, their design and optimization oftentimes require extensive trial-and-errors and case-by-case analysis. Our proposed framework serves as an attempt to mitigate such headaches, offering a formal and systematic approach for analyzing how the accuracy and efficiency of an LLM-based algorithm will be impacted by critical design choices, such as the pattern and granularity of task decomposition, or the prompt for each LLM call. Through a wide range of case studies covering diverse algorithm patterns (including parallel/hierarchical/recursive task decomposition and generic directed acyclic graphs), we demonstrate the proposed framework in action and derive interesting insights that generalize across scenarios, accompanied by systematic empirical validation in synthetic settings.
@@ 1 Introduction
The rapid advancements of pre-trained large language models (LLMs) (Bubeck et al., 2023; Wei et al., 2022a; Schaeffer et al., 2023) have given rise to the new paradigm of algorithmic problem solving, which utilizes LLMs as general-purpose solvers and prompts them with task descriptions and additional inputs that enhance their performance (Wei et al., 2022b; Kojima et al., 2022; Zhang et al., 2024b). However, even the most advanced LLMs today exhibit limitations, such as finite context window sizes and difficulty with complex reasoning. Some of these limitations are fundamental (Merrill & Sabharwal, 2023; Peng et al., 2024; Thomm et al., 2024; Hahn & Rofin, 2024; Wen et al., 2024; Merrill et al., 2024), and resolving them requires new breakthroughs. Moreover, in resource-constrained scenarios where using the state-of-the-art LLMs is not feasible, one might have to resort to smaller and weaker LLMs for solving complex tasks.
@@
All these have motivated the developments of what we call LLM-based algorithms, i.e., algorithms that contain one or multiple LLM calls as sub-routines and fundamentally rely on the capabilities of LLMs. In its most basic form, an LLM-based algorithm can be a combination of one or multiple LLM calls and some prompt engineering (Yao et al., 2023a; Besta et al., 2023; Zhou et al., 2023; Wang et al., 2023). More advanced examples include LLM-powered agentic workflows (Mialon et al., 2023; Pezeshkpour et al., 2024; Xi et al., 2023) and compound AI systems (Zaharia et al., 2024) that augment LLMs with additional abilities like tool use and long-term memory, as well as the emerging paradigm of LLM programming (Schlag et al., 2023; Khattab et al., 2024; Zheng et al., 2024). LLM-based algorithms, in a similar spirit to neuro-symbolic programming (Chaudhuri et al., 2021; Gupta & Kembhavi, 2023) and learning-augmented algorithms (Mitzenmacher & Vassilvitskii, 2022; Lindermayr & Megow, 2022), combine the advantages of LLMs and traditional algorithms. An LLM-based algorithm, designed with human’s intelligence and knowledge of algorithmic problem solving, can exhibit much better controllability and interpretability, stronger performance that is less reliant on delicate prompting, and capabilities that far exceed what a single LLM call can achieve.
Goal and motivations.
A key principle underlying LLM-based algorithms is task decomposition. This is exemplified by the ReAct algorithm (Yao et al., 2023b) that solve a problem via a sequence of reasoning and acting steps, or long-text summarization methods that process the original text by chunks (Kryscinski et al., 2022; Chang et al., 2024). The art of LLM-based algorithm design often lies in figuring out how to decompose the original task into sub-tasks appropriately, so that each can be handled by one LLM call or non-LLM program accurately and efficiently, and the performance of the overall algorithm is guaranteed. Numerous questions can arise in this process, such as:
  • While finer-grained decomposition makes each sub-task easier, it also increases the number of sub-tasks, which might potentially amplify the risk of error accumulation or cascading that could lead to catastrophic failure of the overall algorithm. What would be the appropriate granularity of task decomposition that achieves a balance of these two aspects?
  • Different ways of task decomposition incur different accuracy and efficiency. Does higher accuracy necessarily come with worsened efficiency? When is it possible to achieve the best of both?
  • Long chain-of-thought (CoT) reasoning (Wei et al., 2022b; OpenAI, 2024a; DeepSeek-AI et al., 2025) can enhance the output quality of an LLM call, but also at a substantially higher cost. Is the cost overhead worthy of the gain in accuracy? Should we apply CoT reasoning in solving every sub-task, apply it to a small number of critical sub-tasks only (and tolerate errors in other sub-tasks), or not using CoT at all?
Figure 1: An overview of the proposed analytical framework.
Answering such questions often relies on extensive trial-and-errors in practice. Worse still, the obtained answers are often specific to a particular scenario and do not generalize to others. These challenges call for a more formal and systematic approach of algorithm design and analysis, which might help to predict the empirical performance of LLM-based algorithms, explain curious empirical phenomena, guide the choices of hyperparameters, and potentially inspire better algorithm design.
@@
As an attempt towards achieving this goal, we propose an analytical framework for the design and analysis of LLM-based algorithms, which is summarized in Figure 1 and elaborated in Section 2. The framework represents an LLM-based algorithm as a computational graph (whose nodes correspond to the sub-tasks), and identifies key abstractions such as LLM characteristics (i.e., capabilities and limitations) and LLM inference service. Built upon these, error and cost metrics of the overall algorithm are determined by those of all graph nodes, in line with the task decomposition principle.
We further demonstrate the proposed framework in action for a diverse variety of case studies, deriving interesting and practical insights that generalize across scenarios. In particular, Section 3 investigates parallel decomposition, an elementary yet fundamental pattern of LLM-based algorithms, with a focus on understanding how the granularity of task decomposition impacts the efficiency (Insight 1) and accuracy (Insight 2) of the overall algorithm. Section 4 further studies more general patterns of LLM-based algorithms, offering insights for unified error analysis of generic algorithms under certain technical assumptions (Insight 3), critical design choices for iterative retrieval and reasoning (Insight 4), and making the proposed framework applicable to recursive task decomposition via unrolling (Insight 5).
2 An analytical framework for LLM-based algorithms
This section presents our proposed framework for the design and analysis of LLM-based algorithms.
2.1 Definition and representation
An LLM-based algorithm is simply an algorithm that contains one or multiple LLM calls as its key components. The simplest example would be a single LLM call. More generally, an LLM-based algorithm can be an agentic workflow or compound AI system that utilizes multiple LLM calls and non-LLM programs to handle sub-tasks derived from the original problem, whose outputs together give rise to the final solution of the overall algorithm.
@@ 2.1 Definition and representation
An LLM-based algorithm can be naturally represented as a computational graph. Each graph node takes some inputs from its predecessor nodes, executes certain operations, and returns some outputs. The nodes can be categorized into two types — LLM nodes and non-LLM nodes — as demonstrated in Figure 2. Within an LLM node, the operations consist of a prompter that formats a prompt based on the inputs, an LLM call that processes the prompt [^1] We assume without loss of generality that each LLM node contains one single LLM call, unless specified otherwise., and a parser that extracts the targeted information from the LLM’s response. The prompter and parser, designed for the sub-task of the current node, serve as translators between natural language and traditional data structures. Within a non-LLM node, the operations can be anything that does not involve LLMs. Examples include a symbolic algorithm, an API call for a search engine, or a classical machine learning model.
@@
Given such nodes, the computational graph is built by connecting them with directed edges that specify the data flows within the LLM-based algorithm. For example, Figure 3(a) shows the graph for the pattern of parallel decomposition, which divides the input problem into parallel sub-tasks, solves each of them with one LLM node, and aggregates the results for the final solution; Figure 3(b) illustrates an algorithm for book-length summarization (Chang et al., 2024; OpenAI, 2024b), which divides the input text into multiple chunks and maintains a global summary via incremental updating; and Figure 3(c) represents the ReAct algorithm (Yao et al., 2023b), which consists of multiple iterations of reasoning, tool use and aggregation. The computational-graph formulation is expressive enough to cover these diverse LLM-based algorithms that have been proposed separately in prior works, and facilitates formal analysis for all of them in a unified manner. The computational graph of an LLM-based algorithm can be pre-specified and static, or dynamically constructed at runtime, as will be demonstrated later in our case studies in Sections 3 and 4. Configurations of an LLM-based algorithm include how task decomposition is done (reflected by the graph topology and the sub-tasks of graph nodes), the methods of prompting and parsing for each LLM node, configurations of the LLM(s) being used, among others. The LLM nodes within one graph might use the same or different backbone LLMs.
@@
Figure 2: An LLM node (left) or non-LLM node (right) in the computational graphs of LLM-based algorithms. Each node can have one or multiple inputs/outputs. We use the abbreviation “NL” for “natural language”, and “DS” for “data structure”.
Figure 3: Examples of computational graph representations for LLM-based algorithms. (a) Parallel decomposition. Detailed analysis and concrete examples for this graph pattern will be elaborated in Section 3. (b) Book-length summarization, cf. Figure 1 in (Chang et al., 2024). The “dividing” node contains a symbolic program that divides the input text into multiple smaller chunks. (c) The ReAct algorithm (Yao et al., 2023b). Each “acting” node represents one API call for a certain tool, and each “aggregation” node aggregates the outputs of its predecessor nodes, e.g., by simple concatenation.
2.2 Error and cost analysis
Given an input task and the corresponding LLM-based algorithm, we aim to analyze its performance — how accurately and efficiently the algorithm solves the task — akin to analysis for any generic algorithm. Following the task decomposition principle, error and cost analysis is done by analyzing for each LLM or non-LLM node first, and then for the overall algorithm. Unless otherwise specified, our analysis is deterministic, for a given task instance and fixed random seed(s).
@@ 2.2 Error and cost analysis
Error metrics are task-specific in general. Moreover, the overall algorithm and the LLM nodes within it might have different error metrics. For instance, consider an LLM-based algorithm that answers a given question by first retrieving relevant sentences from the input text with multiple LLM nodes, and then reasoning over the extracted sentences with another LLM node. In this case, the overall algorithm is evaluated by whether it answers the question correctly, while each LLM node responsible for retrieval is evaluated by whether it successfully extracts all necessary sentences within the input text that it handles. [^2] Throughout this work, we use “accuracy” to refer to the broader concept of “quality”, and an “error metric” can be any metric that measures how much the output of an algorithm deviates from certain criteria.
Cost metrics, on the other hand, are largely task-agnostic. Examples include the total length of prompts and total length of generated texts, measured by the numbers of characters or tokens, within one run of the overall algorithm; these metrics are correlated with the monetary costs for using proprietary LLMs via commercial API calls, as well as the time and memory complexities of LLM inference. Another important cost metric is the end-to-end latency of running the overall algorithm; one particular property of this metric is that it can be impacted by parallelism of LLM calls, an important aspect of LLM inference service in practice. Other possible cost metrics include the peak memory usage, the total number of LLM calls, FLOPs, energy consumption, carbon emission, among others. Unless specified otherwise, we focus on the costs of the LLM nodes and neglect those of the non-LLM nodes, since the latter is much smaller than the former in all concrete scenarios that we will consider later in this work.
@@
For error and cost analysis of an LLM-based algorithm, it is necessary to make certain assumptions about the characteristics and inference service of the LLM(s) used within it.
@@
Assumptions on LLM inference service (Yuan et al., 2024; Pope et al., 2023; Zhou et al., 2024), on the other hand, are task-agnostic and only affect the cost metrics, not error metrics. In particular, they determine the relation between the cost metrics and the lengths of prompt and generated text for each LLM call, the parallelism of LLM calls, among others. While LLM inference service in practice can be very diverse, we will see very soon that unified and formal analysis is possible with appropriate abstractions of it.
Error analysis.
The error metrics for the output of each LLM node, which are calculated with respect to what the output should have been if all nodes accomplish their tasks with exact accuracy, depend on the characteristics, i.e., capabilities and limitations, of the LLM, as well as the specific problem instance and random seed(s). For a certain node $v$, the error of its output $y$ can be bounded by some function $f_{v}$ (which depends on the aforementioned factors if $v$ is an LLM node) of the errors of its inputs $x_{1},x_{2},\dots,x_{k}$, i.e., the outputs of its predecessor nodes:
@@
The function $f_{v}$ can be general, for example:
  • $\mathcal{E}(y)\leq\sum_{i\in[k]}\mathcal{E}(x_{i})$ or $\frac{1}{k}\sum_{i\in[k]}\mathcal{E}(x_{i})$. Such linear relationship will appear in a counting task that will be elaborated in Section 3.
  • $\mathcal{E}(y)=\min_{i\in[k]}\mathcal{E}(x_{i})$. In a coding task, the algorithm might generate multiple code samples, and succeed as long as one of the samples is correct (by passing all test cases).
  • $\mathcal{E}(y)\leq\max_{i\in[k]}\mathcal{E}(x_{i})$. This is the case for a sorting task to be introduced in Section 3, where node $v$ is responsible for merging multiple sorted lists and $\mathcal{E}$ stands for the $\ell_{\infty}$ error.
Finally, the error metrics of the overall algorithm are exactly those of the particular graph node that generates the final solution returned by the algorithm.
Cost analysis.
@@
  • Charges of API calls: $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$ are linear functions, namely $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})=c_{\textsf{pre}}\times\mathcal{L}_{\textsf{pre}}$, $\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})=c_{\textsf{dec}}\times\mathcal{L}_{\textsf{dec}}$, where $c_{\textsf{pre}}$ and $c_{\textsf{dec}}$ are the cost per token for the prompt and generated text respectively.
  • Latency: in a memory-bound scenario (Agarwal et al., 2023) where $\mathcal{L}_{\textsf{pre}}$ is limited and the latency of memory IO (for loading model weights and KV caches) is dominant, we have approximately $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})=O(1)$ and $\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})=O(\mathcal{L}_{\textsf{dec}})$; more generally, $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$ might grow linearly or quadratically (e.g., for Transformers with full attention (Vaswani et al., 2017)) with $\mathcal{L}_{\textsf{pre}}$ and $\mathcal{L}_{\textsf{dec}}$.
The cost of the overall algorithm can be written as a function of the cost metrics of all LLM nodes within its computational graph. It might be a simple sum, e.g., for LLM API costs, or a more complex function, e.g., for the end-to-end latency in the presence of parallelism. As an example, consider $k$ independent LLM calls with latencies $\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{k}$ and ideal parallelism of degree $p$, i.e., no extra cost is induced by parallelism; then, the $k$ LLM calls can be divided into $g=\lceil k/p\rceil$ sequential groups, where each group of $p$ parallel LLM calls is bottlenecked by the one with the largest latency, and thus the end-to-end latency becomes
@@
This result will be especially useful for our analysis of parallel decomposition later in Section 3.
Remark 1.
When applying the proposed framework to a new task, we need to first characterize the capabilities and limitations of LLMs for this task (or for the sub-tasks derived from it), e.g., via profiling with a handful of “training samples”. This characterization can be quantitative or qualitative, precise or coarse, depending on the purpose of analysis. For example, in many tasks of interest, a larger problem instance is harder than a smaller one, and thus incurs larger error metrics of an LLM call. For practical purposes like hyperparameter optimization, such coarse characterization might be sufficient already.
Remark 2.